# HF radar estimation of ocean wave parameters: second-order Doppler spectrum versus Bragg wave modulation approach

### v1

@@ Abstract
We propose an original technique for the HF radar estimation of the main sea state parameters by exploiting the amplitude modulation of the radar signal time series. While the classical method for ocean wave measurement is based on the second-order ocean Doppler spectrum, this alternative approach makes use of the dominant first-order Bragg echo in the radar signal, which has a larger signal-to-noise ratio and allows for a farther spatial coverage. We apply this method to a yearly set of HF radar data in the vicinity of Toulon (Mediterranean coast of France) and compare it with the classical second-order method, where we use a nearby buoy record as ground truth. The Doppler spectrum-based method is slightly more accurate in estimating the significant wave height and the mean period at a short distance but the temporal approach allows for a significant increase in range and better captures the variability of wave parameters over time.
journal: Remote Sensing of Environment
@@ 1 Introduction
High-frequency radars (HFR) have been used for about half a century for the monitoring of coastal surface currents (Barrick et al., (1977)). As a secondary geophysical product, HFR also provide an indirect measurement of sea state parameters and wave spectra (e.g. Wyatt, (2021) for a recent review). However, this type of estimation is less accurate and less straightforward than the surface current estimation, for both theoretical and practical reasons. On the theoretical level, while the radial surface current only requires the measurement of the dominant first-order Bragg peaks, the wave spectrum is related to the side-band continuous components of the ocean Doppler spectrum surrounding the Bragg peaks. This second-order contribution has a complex analytical expression in the form of a nonlinear integral equation involving a coupling coefficient and quadratic combinations of the wave spectra at pairs of frequencies satisfying a resonance condition Hasselmann, (1971); Barrick, (1972); Lipa and Barrick, (1986). Under some restrictive assumptions, Barrick, 1977b ; Barrick, 1977a derived in $1977$ a method to extract the main sea state parameters from this second-order integral and accordingly, from the measured ocean Doppler spectrum. He showed that the significant wave height ($H_{s}$) and the peak wave period ($T_{p}$) can be expressed as simple ratios of weighted integrals involving the first- and second-order Doppler spectra. Soon after, linear and nonlinear inversion methods for retrieving the ocean wave spectrum from one single radar site were developed in a series of papers (e.g. Lipa, (1977); Wyatt, (1990); Hisaki, (1996); Wyatt, (2021)). However, the success of these methods for inverting the sea state parameters or the wave spectrum is bound to the ability to measure the second-order Doppler spectrum with a sufficient signal-to-noise ratio. Except for the high sea state and the highest radar frequencies, the second-order Doppler spectrum is usually several tens of dB below the level of the first-order peaks and reaches sooner the noise threshold when the range is increased. Hence, wave estimation methods based on the second-order Doppler analysis are confined to a restricted radar coverage in the near to middle range, depending on the radar frequency. Another issue is the clear identification of the second-order contribution and its separation from the first-order Bragg components, which can be ambiguous especially in the case of strong currents or in the absence of azimuthal resolution. For these reasons, empirical methods using the second-order Doppler spectrum (e.g. Essen et al., (1999); Gurgel et al., (2006); Lopez et al., (2016)) or empirical adjustments of Barrick’s method Heron and Heron, (1998); Maresca Jr and Georges, (1980); Ramos et al., (2009); Alattabi et al., (2019); Al-Attabi et al., (2021); Basañez et al., (2020) are often preferred to the original approach of Barrick. Note that these methods necessitate the calculation of Doppler spectra that are resolved both in range and azimuth and can thus be associated with a reduced sea surface patch with constant sea state parameters. In the case of compact antenna arrays, the Doppler spectra are omnidirectional, that is, they cannot be resolved in azimuth. This does not allow for the measurement of the wave frequency spectrum from a single station. Nevertheless, the $H_{s}$ can be estimated on range arcs adapting Barrick’s original methods. The long-term performances of such measurements have been assessed by comparison with buoys and models and are quite variable depending on sea state, range, and direction of waves (see e.g. Roarty et al., (2015); Orasi et al., (2018); Saviano et al., (2020) or Lorente et al., (2022) for a recent review).
@@
Figure 1: Location of HF radar coverage a) in the western Mediterranean Sea and b) zooming in on the area. The red dot marks the location of Fort Peyras, the blue dot that of the CANDHIS buoy, and the white lines represent the resolution in ranges (every $7.5km$) and azimuths (every $5\rm{{}^{o}}$) of the radar coverage. The orange line a40 marks the main azimuth and the orange arcs of circles mark the range numbers r8, r20, and r35 used for the HFR comparison. The background color is the bathymetry measured in $m$. Toponyms cited in the main text appear in the figure.
@@
In this paper, we propose a novel approach to sea state estimation based on the analysis of the first-order Bragg component of the Doppler spectrum. We exploit the amplitude modulation of the received radar time series to infer the $H_{s}$ and the $T_{p}$. The idea of using the first-order backscattered signal to estimate the sea state parameters has been recently put forward in Shahidi and Gill, (2020); Shahidi et al., (2024) based on some mathematical approximations starting from the expression of the electric field backscattered from a patch of the sea surface. It has been found that the $H_{s}$ is simply proportional to the variance over time of the first-order received electric field and that the $T_{p}$ can be obtained from the peak frequency of the same times series after integration over range. The method has been assessed with limited but promising experimental tests. Our approach, which was developed independently, is based on simple physical heuristic arguments and reaches the same conclusion that the sea state parameters can be derived from the statistical properties of the radar signal time series (in which the first-order Bragg contribution is dominant) and does not require any Doppler processing nor the identification of the second-order contribution. Our formulation, however, is quite different in that it relies on the calculation of the analytical and anti-analytical radar signal, a technique that allows us to isolate the contribution of long waves in the backscattered field.
@@
The paper serves a double objective. First, we test the performance of Barrick’s method based on a yearly data set of HF measurements with a radar located near Toulon. Its accuracy is very sensitive to some tuning parameters in its numerical implementation, which is part of the tacit know-how of the users and is rarely mentioned in the literature. A key step is the identification of the separation between the first- and second-order Doppler spectrum, for which we propose an empirical but explicit criterion. The second objective is to test the aforementioned new method based on the temporal properties of the backscattered signal. The HFR data set is described in section 2, together with in-situ measurements and numerical models which are used for the assessment. The radar processing methods are presented in Section 3 and a variety of statistical comparisons over one year of data is given in Section 4 while discussing the results obtained.
Figure 2: a) Mean value (in meter) of $H_{s}$ over the year 2020 in the radar coverage according to the WWIII model; b) Root-mean-square difference (in meter) between the $H_{s}$ at any location and the $H_{s}$ at the CANDHIS buoy location, according to the WWIII model. The black dashed line marks the azimuth 40 and the black circles mark the ranges r8, r12, and r35, respectively, used for the HFR comparison (see Figure 1).
@@ 2.1 High Frequency Radar data
In this study, we use both omnidirectional and directional monostatic signals collected by the Fort Peyras site (Figure 1) during $2020$, a year for which data availability was over $91\%$. This radar operates in a $16.15$ MHz center frequency band with an assigned bandwidth of $100$ kHz. Depending on weather conditions, it can cover a range of $60$ km with a range resolution of $1.5$ km and an angular sector of about $110\rm{{}^{o}}$. The data consist of the complex I/Q voltage recorded on the set of $12$ receive antennas and sorted in range after a preliminary Doppler processing in the fast time. The sampling rate in the slow time is $0.26$ seconds and the time series are acquired in continuous blocks of $4096$ values (lasting about $18$ min) every $20$ min, which corresponds to the standard “SORT file” output format of WERA HF radar. The data are further processed in azimuth by applying the classical Beam Forming algorithm to the $12$ channels of the antenna array. This leads to a complex discrete time series $s(t)$ for each range and azimuth and each $20$ min sequence. Not applying the Beam Forming technique results in an omnidirectional signal for each antenna channel, with which it is only possible to know from what distance the signal is being received but not from which bearing.
@@ 2.2 Wave data
For the year $2020$ the time series of $H_{s}$ and $T_{p}$ are collected by a CANDHIS (Centre d’Archiveage National de Données de Houle In-Situ) buoy located in the vicinity of Porquerolles island ($42\rm{{}^{o}}58.00^{\prime}$, $6\rm{{}^{o}}12.29^{\prime}$, see Fig. 1). The data buoy are provided by the CEREMA (Climat et Territoires de Demain) and they are available in Cerema, (2024). The time resolution of the buoy data is $30$ minutes. During the same year, the results of the Wavewatch III model (WWIII) were available over the study area on an unstructured grid (with a spatial resolution between $400$m near the coast and $25$km offshore) forced by $10$-m wind fields obtained from the non-hydrostatic mesoscale Weather Research and Forecast model (for more details consult Lira-Loarca et al.,, 2022). The WWIII model has an hourly time resolution.
While the ground truth in the vicinity of the radar area is limited to the single location of the CANDHIS buoy, it is possible to use the WWIII model to complement the analysis. The spatial variability of $H_{s}$ under the radar coverage has been evaluated by calculating the root mean square difference over the whole year 2020 between the modeled $H_{s}$ at any location in the Toulon area and the modeled $H_{s}$ at the CANDHIS buoy location (Figure 2). As can be seen, this difference ranges from about $10$ cm at short distances from the radar to about 40 cm at long distances. Accepting this small to moderate difference, we can use the unique buoy measurement point as ground truth for validating the wave parameters estimated with the HFR dataset.
@@ 3 Material and Methods
Two different methods are used to estimate the ocean waves, $H_{s}$ and $T_{p}$, from HF radar signals. The first is the classical Second Order Doppler Spectrum (SODS) method, first introduced by Barrick, 1977a ; Barrick, 1977b , while the second is an original method that uses only the temporal variability of the first order peak and will be referred to as the “First Order Bragg Peak” (FOBP) method.
@@ 3.1 Second Order Doppler Spectra Method
The classical method for estimating $H_{s}$ and $T_{p}$ from the HF sea echo originates from the seminal work of Barrick, 1977a ; Barrick, 1977b and involves straightforward calculations based on the ocean Doppler spectrum. The hourly range-azimuth resolved Doppler spectra are classically obtained by forming the power spectrum of the radar signal:
$\sigma(\omega)=\langle\lvert\hat{s}(\omega)\rvert^{2}\rangle,$ (1)
where $\hat{s}(\omega)$ is computed in the standard way on a discrete set of frequencies using a Fast Fourier Transform of the time series with a Hamming attenuation window and an integration time of about $2$ min. To reduce the noise variance, a further incoherent summation is applied by taking a median average of the sample spectra $\lvert\hat{s}(\omega)\rvert^{2}$ over $15$ overlapping intervals within the $18$ min duration of each available SORT file (3 per hour). This will result in an average Doppler spectrum with a bandwidth of $4$ Hz (corresponding to a sampling rate of $0.26$ sec) and a frequency resolution of $\Delta f=1/T\simeq 0.008$ Hz. An example of such a Doppler spectrum is shown in Figure 3, a) for the $8$ range cell ($12$ km) and azimuth $40$ (marked with orange lines in Figure 1, b) for March $8$, $2020$. The corresponding range Doppler spectrum along the same direction is shown in Figure 4. In this example, the strongest first-order Bragg line extends to about $90$ km, while the strongest second-order sideband is visible up to $65$ km away from the radar. The dashed white circles highlight a power grid artifact ($50$ Hz harmonics) that contaminates the low Doppler frequency content of the radar signal at some ranges; this electrical artifact is particularly pronounced at the Fort Peyras site due to the imperfect power supply at this location. It does not affect the estimation of the radial current from the first-order Bragg lines but must be treated with caution when dealing with the full ocean Doppler spectrum.
@@
$\omega_{B}=2\pi\sqrt{\frac{g}{\pi\lambda}}$ (3)
is the Bragg frequency and $g=9.81\ m.s^{-2}$ is the gravitational constant.
The $T_{p}$ can also be determined from a different ratio of integrals based solely on the second-order Doppler spectrum:
@@
where the bounds of the integrals can be either 00 and $\omega_{B}$ or $\omega_{B}$ and +$\infty$+ ∞, assuming that the positive frequency sidebands are the strongest (otherwise one should take the symmetric frequency intervals on the negative side). In this study, the maximum value of the integral of both intervals (from 00 to $\omega_{B}$ and from $\omega_{B}$ to +$\infty$+ ∞) is chosen for each time. The main challenge of this method is to effectively discriminate between first and second-order spectra and to accurately distinguish signal from noise. To overcome this, specific techniques have been developed that allow the precise setting of separation thresholds and the establishment of control criteria for integration regions. The effectiveness of this technique is limited by its dependence on the second-order Doppler spectrum, which restricts its application to shorter distances where this spectrum remains well above the noise level.
To determine the exact point of transition from first to second-order spectra, and thus avoid false boundaries, the following method is used: positive peaks (marked with a blue circle in Figure 3, b) and negative peaks (marked with red circles in Figure 3, b) are selected. The minimum negative peak (marked with a solid red circle in Figure 3, b) and the maximum positive peak at frequencies higher than the minimum negative peak (marked with a solid blue circle in Figure 3, b) are identified. We impose the condition that the intensity difference between the first-order peak and the minimum negative peak (d1) must be at least twice the difference between the minimum negative peak and the maximum positive peak (d2). If this condition is not met, the selected minimum negative peak is discarded and the process is repeated with another negative peak initially identified. At the same moment, the noise level is determined as the average intensity for frequencies higher than $1.75$ Hz (see Figure 3, a).
@@
3.2 First-order Bragg peak method
To achieve wave estimation over longer distances, a new method is proposed where working with the second-order spectrum becomes impossible. Instead, we exploit the modulation of the first-order Bragg signal by long waves, defined as waves significantly longer than the Bragg wave ($\omega_{p}<<\omega_{B}$). In a two-scale image, Bragg waves ride on long waves whose orbital motion modulates their phase velocity and amplitude. This process comes into play through a non-linear interaction between short and long waves, which heuristically involves the product of these two components. Assuming that the first-order Bragg scattering mechanism is dominant in the temporal radar signal $s(t)$, and considering both the approaching (positive frequencies) and receding (negative frequencies) components of the long waves and the Bragg waves, we can postulate the following simplified form for the received complex time series:
$s(t)\sim(R^{+}e^{i\omega_{p}t}+R^{-}e^{-i\omega_{p}t})(a^{+}e^{i\omega_{B}t}+a^{-}e^{-i\omega_{B}t})e^{i\omega_{u}t}$ (5)

### v2

@@ Abstract
We propose an original technique for the HF radar estimation of the main sea state parameters by exploiting the amplitude modulation of the radar signal time series. While the classical method for ocean wave measurement is based on the second-order ocean Doppler spectrum, this alternative approach uses the slow amplitude modulation of the Bragg wave in the radar signal, which is more robust to noise than the second-order echo. We apply this method to an annual set of HF radar data in the vicinity of Toulon (Mediterranean coast of France) and compare it with the classical second-order method, using a nearby buoy and the WWIII model as ground truth. The Doppler spectrum-based method is found to be more accurate in calculating the significant wave height and the peak wave frequency while the modulation approach provides coarser estimates but can achieve longer ranges and higher temporal coverage.
journal: Coastal Engineering
@@ 1 Introduction
Figure 1: Location of HF radar coverage a) in the western Mediterranean Sea and b) zooming in on the area. The red dot marks the location of Fort Peyras, the blue dot that of the CANDHIS buoy, and the white lines represent 5 range units (every $7.5km$) and 5 azimuth units (every $5\rm{{}^{o}}$) of the radar coverage. The background color is the bathymetry measured in $m$. Toponyms cited in the main text appear in the figure.
High-frequency radars (HFR) have been used for about half a century for the monitoring of coastal surface currents (Barrick et al., (1977)). As a secondary geophysical product, HFR also provide measurement of sea state parameters and wave spectra (e.g. Wyatt, (2021) for a recent review). However, this type of estimation is less accurate and less straightforward than the surface current estimation, due to theoretical as well as practical reasons. On the theoretical level, while the radial surface current only requires the measurement of the dominant first-order Bragg peaks, the wave spectrum is related to the side-band continuous components of the ocean Doppler spectrum surrounding the Bragg peaks. This second-order contribution has a complex analytical expression in the form of a nonlinear integral equation involving a coupling coefficient and quadratic combinations of the wave spectra at pairs of frequencies satisfying a resonance condition Hasselmann, (1971); Barrick, (1972); Lipa and Barrick, (1986). Under some restrictive assumptions, Barrick, 1977b ; Barrick, 1977a derived in $1977$ a method to extract the main sea state parameters from this second-order integral and accordingly, from the measured ocean Doppler spectrum. He showed that the significant wave height (SWH) and the centroïd wave frequency can be expressed as simple ratios of weighted integrals involving the first- and second-order Doppler spectra. Soon after, linear and nonlinear inversion methods for retrieving the ocean wave spectrum from single or dual radar sites were developed in a series of papers (e.g. Lipa, (1977); Wyatt, (1990); Hisaki, (1996)). Today, the inversion of the waves has reached a mature state and there are well-established techniques and commercial software for the determination of the ocean wave directional spectrum (see e.g. Wyatt, (2021) for a recent review). However, these inversion techniques require more expertise than that required to calculate the main wave parameters using Barrick’s formulas. They have also physical as well as practical limitations (Wyatt, (2000)) that restrict their range of applicability. The most critical issue relates to the quality of the radar data and, in particular, to the ability to measure the second-order Doppler spectrum with a sufficient signal-to-noise ratio. Except for the high sea state and the highest radar frequencies, the second-order Doppler spectrum is usually several tens of dB below the level of the first-order peaks and reaches sooner the noise threshold when the range is increased. Hence, wave estimation methods based on the second-order Doppler analysis are confined to a restricted radar coverage in the near to middle range, depending on the radar frequency. Another issue is the clear identification of the second-order contribution and its separation from the first-order Bragg components, which can be ambiguous especially in the case of strong currents or in the absence of azimuthal resolution. For these reasons, many authors have only considered the estimation of the bulk wave parameters using empirical methods (e.g. Essen et al., (1999); Gurgel et al., (2006); Lopez et al., (2016)) or empirical adjustments of Barrick’s method Heron and Heron, 1998a ; Maresca Jr and Georges, (1980); Ramos et al., (2009); Alattabi et al., (2019); Al-Attabi et al., (2021); Basañez et al., (2020). Most of these methods necessitate the calculation of Doppler spectra that are resolved both in range and azimuth and can thus be associated with a reduced sea surface patch with constant sea state parameters. In the case of compact antenna arrays, the Doppler spectra are nondirectional, that is, they cannot be resolved in azimuth. This does not allow for the measurement of the wave frequency spectrum from a single station. Nevertheless, the $H_{s}$ can be estimated on range arcs adapting Barrick’s original methods. The long-term performances of such measurements have been assessed by comparison with buoys and models and are quite variable depending on sea state, range, and direction of waves (see e.g. Roarty et al., (2015); Orasi et al., (2018); Saviano et al., (2020) or Lorente et al., (2022) for a recent review).
@@
In this paper, we propose a novel approach to sea state estimation based on the time variation of the total radar signal, without applying the classical separation between first- and second-order components; instead, we exploit the amplitude modulation of the received voltage to infer the SWH and the peak wave frequency. The idea of using the time variation of the radar signal and bypassing the Doppler spectrum to estimate the sea state parameters has been recently proposed in Shahidi and Gill, (2020); Shahidi et al., (2024) based on some mathematical approximations starting from the expression of the first-order electric field backscattered from a patch of the sea surface. It has been found that the SWH is simply proportional to the variance of the received voltage and that the peak frequency can be obtained from the same time series after integration over range. The method has been assessed with limited but promising experimental tests. Our approach, developed independently, is based on simple physical heuristic arguments and reaches a similar conclusion that the sea state parameters can be derived from the statistical properties of the received voltage and does not require any Doppler processing or identification of the second order contribution. However, our formulation is quite different in that it does not reduce the signal to the sole contribution of the first-order Bragg term, but also includes higher-order terms through the hydrodynamic modulation of the backscattering. In addition, it is based on the calculation of the analytical and anti-analytical radar signal, a technique that allows us to isolate the contribution of long waves in the backscattered field.
@@
The paper serves a double objective. First, we test the performance of Barrick’s method based on an annual data set of HF measurements with a radar located near Toulon. Its accuracy is very sensitive to some tuning parameters in its numerical implementation, which is part of the tacit know-how of the users and is rarely mentioned in the literature. A key step is the identification of the separation between the first- and second-order Doppler spectrum, for which we propose an empirical but explicit criterion. The second objective is to test the aforementioned new method based on the temporal properties of the backscattered signal. The HFR data set is described in section 2, together with in-situ measurements and numerical models which are used for the assessment. The radar processing methods are presented in Section 3 and a variety of statistical comparisons over one year of data is given in Section 4 while discussing the results obtained.
Figure 2: Annual mean value of a) the SWH (in meter) ; b) the peak wave frequency in the radar coverage according to the WWIII model. The blue dot marks the position of the CANDHIS buoy.
@@ 2.1 High Frequency Radar data
In this study, we use the monostatic signals collected by the Fort Peyras site (Figure 1) during $2020$, a year for which data availability was over $91\%$. This radar operates in a $16.15$ MHz center frequency band with an assigned bandwidth of $100$ kHz. Depending on weather conditions, it can cover a range of $60$ km with a range resolution of $1.5$ km and an angular sector of about $110\rm{{}^{o}}$. The data consist of the complex I/Q voltage recorded on the set of $12$ receive antennas and sorted in range after a preliminary Doppler processing in the fast time. The sampling rate in the slow time is $0.26$ seconds and the time series are acquired in continuous blocks of $4096$ values (lasting about $18$ min) every $20$ min, which corresponds to the standard “SORT file” output format of WERA HF radar. The data are further processed in azimuth by applying the classical beamforming algorithm (with steering steps of one degree) to the $12$ channels of the antenna array. This leads to a complex discrete time series $s(t)$ for each radar cell of size $1.5$ km×$\times 1$ degree and each $20$ min sequence. In the following, each radar cell is labeled by a range number ($1$ to $64$ starting from the radar) and an azimuth number ($1$ to $110$ starting from the westernmost corner).
@@ 2.2 Wave data
We recall the standard definitions of the SWH as the integral of the wave frequency spectrum $S(f)$:
[…]
which can also be obtained as four times the RMS elevation of the free surface. The peak or dominant frequency is defined as:
$f_{p}=\mathrm{ArgMax}S(f)$ (2)
and the mean or centroïd wave frequency $f_{c}$ is defined with the first moment of the wave frequency spectrum $S(f)$:
[…]
As we will see, the peak frequency $f_{p}$ turns out to be more relevant than the centroïd wave frequency ${f_{c}}$ for the comparisons with radar-based estimates, at least in the case of a wind-wave sea with non distinct swell component, as it is in general the case in the area of Toulon.
For the year $2020$ we used as ground truth a time series of wave parameters collected by a CANDHIS (Centre d’Archiveage National de Données de Houle In-Situ) buoy located in the vicinity of Porquerolles island, at about $32$ km from the radar ($42\rm{{}^{o}}58.00^{\prime}$, $6\rm{{}^{o}}12.29^{\prime}$, see Fig. 1). These data are provided by the CEREMA (Climat et Territoires de Demain) at a temporal resolution of $30$ minutes and they are available in Cerema, (2024). They cover the whole year 2020 with the exception of a 83 days gaps between May, 5 and July, 27. They provide the $H_{s}$ calculated with four times the RMS elevation and the peak frequency $f_{p}$.
During the same year, the results of the Wavewatch III model (WWIII) at an hourly time resolution were available over the study area on a 3-level nested grid (with a spatial resolution of $800$m in the radar aera) forced by $10$-m wind fields obtained from the non-hydrostatic mesoscale Weather Research and Forecast model (for more details consult Lira-Loarca et al.,, 2022). The SWH and the mean wave frequency are estimated through the spectral moments (1) and the peak wave frequency $f_{p}$ is also provided. The annual mean values of the SWH and the peak wave frequency in the coastal region of Toulon according to the WWIII model are shown in Figure 2. As seen, the typical SWH in the radar aera is of the order of $1$ m while the typical value of the peak frequency is about $0.22$ s. While the ground truth in the vicinity of the radar area is limited to the single location of the CANDHIS buoy, it is possible to use the WWIII model to complement the analysis. To check the accuracy of the WWIII in the radar area, we compared the time series of $H_{s}$ and $f_{p}$ at the buoy location to the actual measurements. The point by point comparison is shown in the scatterplots of Figures 3-4. The Root Mean Square Difference (RMSD) of SWH is $25$ cm, which is about $25\%$ of its mean value ($0.96$ m) while the RMSD of the peak wave frequency is $0.037$, which is about $20\%$ of its mean value ($0.19$ Hz). The Pearson correlation coefficient is higher for the SWH ($0.94$) than for the peak wave frequency ($0.70$). The statistical parameters to compare the different time series are summarized in Table 1.
Figure 3: Time series (top) and scatterplot (bottom) of the SWH (in meter) over the year 2020 at the CANDHIS location according to the WWIII model (black line) and according to the buoy (red lines).
Figure 4: Same as Figure 3 for the peak wave frequency
Table 1: Mean wave parameters from the CANDHIS buoy and estimation of the WWIII/CANDHIS difference over the year $2020$.
CANDHIS WWIII/CANDHIS
mean RMS mean bias RMSD R
$H_{s}(m)$ $0.96$ $0.71$ $-0.03$ $0.25$ $0.94$
$f_{p}$ (Hz) $0.19$ $0.044$ $0.004$ $0.037$ $0.70$
@@ 3 Material and Methods
Two different methods have been used to estimate the ocean waves, $H_{s}$ and $f_{p}$, from HF radar signals. The first is the classical Second Order Doppler Spectrum method, first introduced by Barrick, 1977a ; Barrick, 1977b , while the second is an original method that uses only the time modulation of Bragg wave amplitude. In the following we will refer to these methods simply as “Barrick’s method” and the “Modulation method”.
@@ 3.1 Second Order Doppler Spectra Method
The classical method for estimating $H_{s}$ and $f_{p}$ from the HF sea echo originates from the seminal work of Barrick, 1977a ; Barrick, 1977b and involves straightforward calculations based on the ocean Doppler spectrum. The hourly range-azimuth resolved Doppler spectra are classically obtained by forming the power spectrum of the radar signal:
$\sigma(f)=\langle\lvert\hat{s}(f)\rvert^{2}\rangle,$ (4)
where $\hat{s}(f)$ is computed in the standard way on a discrete set of frequencies using a Fast Fourier Transform of the time series with a Hamming attenuation window and an integration time of about $2$ min. To reduce the noise variance, a further incoherent summation is applied by taking a median average of the sample spectra $\lvert\hat{s}(f)\rvert^{2}$ over $15$ overlapping intervals within the $18$ min duration of each available SORT file (3 per hour). This will result in an average Doppler spectrum with a bandwidth of $4$ Hz (corresponding to a sampling rate of $0.26$ sec) and a frequency resolution of $\Delta f=1/T\simeq 0.008$ Hz. An example of such a Doppler spectrum is shown in Figure 5, a) for the $8$ range cell ($12$ km) and azimuth $40$ (marked with orange lines in Figure 1, b) for March $8$, $2020$. The corresponding range Doppler spectrum along the same direction is shown in Figure 6. In this example, the strongest first-order Bragg line extends to about $90$ km, while the strongest second-order sideband is visible up to $65$ km away from the radar. The dashed white circles highlight a power grid artifact ($50$ Hz harmonics) that contaminates the low Doppler frequency content of the radar signal at some ranges; this electrical artifact is particularly pronounced at the Fort Peyras site due to the imperfect power supply at this location. It does not affect the estimation of the radial current from the first-order Bragg lines but must be treated with caution when dealing with the full ocean Doppler spectrum.
@@
$f_{B}=\sqrt{\frac{g}{\pi\lambda}}$ (6)
is the Bragg frequency and $g=9.81\ m.s^{-2}$ is the gravitational constant. The mean wave frequency can also be determined, in principles, from a different ratio of integrals based solely on the second-order Doppler spectrum:
@@
where the bounds of the integrals can be either 00 and $f_{B}$ or $f_{B}$ and +$\infty$+ ∞, assuming that the positive frequency sidebands are the strongest (otherwise one should take the symmetric frequency intervals on the negative side). In this study, the maximum value of the integral of both intervals (from 00 to $f_{B}$ and from $f_{B}$ to +$\infty$+ ∞) is chosen for each time. In reality, the maximum available Doppler frequency for the second-order Doppler spectrum extends at best to 3 or 4 times the Bragg frequency (i.e. about $1.5$ Hz at our operating radar frequency $16.15$ MHz), so that the above integral is actually closer to the peak frequency than to the mean wave frequency.
The main challenge of Barrick’s method is to effectively discriminate between first and second-order spectra and to accurately distinguish signal from noise. To overcome this, specific techniques have been developed that allow the precise setting of separation thresholds and the establishment of control criteria for integration regions. The effectiveness of this technique is limited by its dependence on the second-order Doppler spectrum, which restricts its application to shorter distances where this spectrum remains well above the noise level. Due to this greater sensitivity to noise, it is important to subtract the mean noise level from the second-order Doppler spectrum before performing the integration in (5) and (7).
To determine the exact point of transition from first to second-order spectra, and thus avoid false boundaries, the following method is used: positive peaks (marked with a blue circle in Figure 5, b) and negative peaks (marked with red circles in Figure 5, b) are selected. The minimum negative peak (marked with a solid red circle in Figure 5, b) and the maximum positive peak at frequencies higher than the minimum negative peak (marked with a solid blue circle in Figure 5, b) are identified. We impose the condition that the intensity difference between the first-order peak and the minimum negative peak ($d_{1}$) must be at least twice the difference between the minimum negative peak and the maximum positive peak ($d_{2}$). If this condition is not met, the selected minimum negative peak is discarded and the process is repeated with another negative peak initially identified. This threshold value of $2$ is purely empirical and has been found to be effective in removing satellite peaks near the first-order Bragg peak that could be misinterpreted as second-order peaks. Although these peaks are located at very close frequencies, the resulting error in calculating the integral of the second-order Doppler spectrum can be large, resulting in an inaccurate estimate of $H_{s}$.
At the same moment, the noise level is determined as the average intensity for frequencies higher than $1.75$ Hz (see Figure 5, a). In spite of these tests, the correct identification of the second-order Doppler spectrum may sometimes fail and lead to erroneous estimates of the wave parameters. However, a significant proportion of these irrelevant values can be eliminated by a posteriori quality checks such as the removal of outliers in the time series of estimated wave parameters.
@@
3.2 The Bragg Wave Modulation Method
Figure 7: a) Real part of the received radar signal complex time series (black solid line) for the range $10$ ($15$ km) and azimuth $75$ on September $27$, $2020$, at $23$h where $H_{s}$ and $T_{p}$ from CANDHIS are $2.93\ m$ m and $7.5\ s$, respectively. The red solid line is the signal envelope, obtained with the modulus (and its opposite value) of the analytical signal. b) Modulus of the analytical signal for the same time. c) Sum of the analytical and anti-analytical power spectra ($\sigma_{\lvert s_{a}^{+}\rvert}(f)+\sigma_{\lvert s_{a}^{-}\rvert}(f)$) for the same time. The common wave frequencies in the Mediterranean Sea are marked with a red rectangle where the maximum power spectrum is related to peak wave frequency ($f_{p}$)
It is well known that the backscatter from short waves is modulated by the slow variation of long waves through nonlinear hydrodynamical interaction mechanisms. The variation of the received power from the resonant short wave is usually represented by the so-called Modulation Transfer Function (MTF). Even though the concept of MTF has been mostly defined in the context of microwave scattering (e.g. Keller and Wright, (1975); Keller et al., (1994)), it also applies to HFR whenever there are waves in the ocean that are much longer than the Bragg wave. Here we adapt the concept of MTF to the simplified cased where there is only one long wave (say, the peak wave). Its frequency and wavenumber are assumed to be much smaller than those of the Bragg frequency ($f_{p}<<f_{B}$, $k_{p}<<k_{B}$) while its amplitude is much larger ($A_{p}>>A_{B}$) so we can assume a simple Two-Scale model. Considering the hydrodynamic interaction of the lowest order in long-wave steepness, we have that both the amplitude and wavenumber of the Bragg wave are modulated by the long-wave orbital cycle $\Phi$ (Longuet-Higgins and Stewart,, 1960):
$\tilde{a}_{B}(t)=a_{B}(1+k_{p}a_{p}\cos\Phi)$ (8)
$\tilde{k}_{B}(t)=k_{B}(1+k_{p}a_{p}\cos\Phi),$ (9)
Assuming that the dominant part of the sea echo is due to first-order Bragg scattering and allowing the Bragg wave to propagate in both directions, a similar amplitude modulation will result for the complex radar signal:
$s(t)\sim A_{B}^{+}(t)e^{2{\rm{i}}\pi(\tilde{f}_{B}+f_{u})t}+A_{B}^{-}(t)e^{-2{\rm{i}}\pi(\tilde{f}_{B}-f_{u})t}$ (10)